OpenStudy (anonymous):

write an equation of the line containing the given points. A(2,7), B(3,4)

OpenStudy (anonymous):

use (y2-y1)=m(x2-x1) where your points are (x1,y1) and (x2,y2) that will give you m (the slope) then use (y-y1)=m(x-x1) (or you can use x2, y2 if you choose) and then solve for y

OpenStudy (anonymous):

ok whats the second formula called

OpenStudy (anonymous):

they're both called point-slope formula. they're the same formula essentially, you're just pluggin in different things you know to figure out what you don't know. like, for the first part you know two points but not the slope, so you have to find that. and in the second part you want the whole equation of the line, but you only need 1 point and the slope.

OpenStudy (anonymous):

oh okay thank you